OpenStudy (arbyj):

Someone help me solve; 5(3y+2)=13+3y (please show work)

OpenStudy (anonymous):

Expand the right side of the equation.

OpenStudy (arbyj):

15y+10=13+3y

OpenStudy (arbyj):

@bloopman

OpenStudy (anonymous):

Subtract 3y from both sides of the equation.

OpenStudy (anonymous):

so you have: 5(3y+2) = 13+3y <<<First simplify the mess... distribute and combine like terms 15y+10 = 13+3y <<< get all of the same term on one side (start with the variables) 12y = 3 <<<divide to solve for the variable y = 1/4

OpenStudy (arbyj):

12y+10=13 @bloopman

OpenStudy (anonymous):

Alright, finally, subtract 10 from both sides. 12y = 3. Divide both sides by 12 and simplify.

OpenStudy (arbyj):

y=.25 thank you! :)